What is the work function of gold metal in kJ/mol if light with = 234 nm is necessary to eject electrons?
1
Expert's answer
2019-09-16T04:06:07-0400

E = hc/λ

E= (6.626 x 10^-34 x 3 x 10^ 8) 1x10^9/234=2.83162393x10^-19 J= 2.83162393x 10^-22

2.83162393x 10^-22 x 6.02 x 10^23 = 170.463761 kJ/mol